OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 0}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\quad \textit{point-slope form, plug in your values}\)

OpenStudy (anonymous):

@jdoe0001 So, y - 4 = 6/-2 (x - 4) ?

OpenStudy (anonymous):

And then I would distribute right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

hmmm notice your values though -> y - 4 = 6/-2 (x - 4) ^ ^ (4, 4) ?

OpenStudy (jdoe0001):

you have (4,0) and (2,6) to use, you don't have a (4,4) point

OpenStudy (jdoe0001):

so using say (4, 0) will be y - 0 = 6/-2 (x - 4) ^ ^

OpenStudy (jdoe0001):

then distribute